OpenStudy (anonymous):

What is the equation of the line whose y-intercept is 3 and slope is 1? y = x - 3 y = x + 3 y = 3x + 1

OpenStudy (4everaddicted2anime):

y=x+3

OpenStudy (anonymous):

thanks

OpenStudy (sushi121212):

It's y=x+3 because the template would be: y=mx+b, in which b is the y intercept and the slope is the m. then, just plug it in: y=x+3 (you don't have to plug in the 1 for the slope)